# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ III Proof Theorem 2
00 0 $=\nabla F({\bf{X}}^{\star})+{\bf{W}}_{b}{\bf{U}}^{\star},$ (44a)
00 0 $={\bf{W}}_{b}{\bf{Z}}^{\star},$ (44b)
@@
$\mathcal{E}_{\mathrm{s}}^{t}\triangleq\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]\text{ and }\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
@@
Proof See Appendix XII. $\Box$ With this error dynamics, similar as Lemma 3, we give the following descent inequalities.
@@
where $\vartheta_{\mathrm{s}}=\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}$, $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3(1-p)(2+p^{2})}{\chi^{2}}$
@@
Proof See Appendix XIII. $\Box$
@@
$\chi>\max\left\{\frac{1}{p},\frac{36}{1-\lambda_{2}},\frac{72(1-p)}{1-\lambda_{2}}\right\},$ (48)
@@
Proof See Appendix XIV. $\Box$
IV Proof of Lemma 2
Proof It follows from (2b), $\mathbf{I}-\mathbf{W}_{a}=\frac{1}{2\chi}\mathbf{W}_{b}^{2}$ and ${\bf{E}}^{t}=\frac{1}{2\chi}(\theta_{t}-1){\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$ that
@@ IV Proof of Lemma 2
$={\bf{W}}_{a}\hat{\mathbf{Z}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t}.$
Since $\beta=1$, it follows from (2b), (2c), and $\mathbf{I}-\mathbf{W}_{a}=\frac{1}{2\chi}\mathbf{W}_{b}^{2}$ that
${\bf{Y}}^{t+1}={\bf{Y}}^{t}+\frac{1}{2\chi}{\bf{W}}_{b}^{2}\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}={\bf{Y}}^{t}+({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}.$
By ${\bf{R}}^{t}={\bf{Y}}^{t}+\alpha\nabla F(\bar{{\bf{X}}}^{t})$, ${\bf{\Sigma}}_{2}^{t}=\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1})$, and ${\bf{E}}^{t}=\frac{1}{2\chi}(\theta_{t}-1){\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, we have
@@
$=({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}+\alpha(\nabla F(\bar{{\bf{X}}}^{t+1})-\nabla F(\bar{{\bf{X}}}^{t}))$
$=({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}.$
@@
${\bf{X}}^{t+1}$ $={\bf{W}}_{a}\hat{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$
${\bf{R}}^{t+1}$ $={\bf{R}}^{t}+({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}+{\bf{W}}_{b}{\bf{E}}^{t},$
which also can be rewritten as (since ${\bf{W}}_{a}={\bf{I}}-\frac{1}{2\chi}{\bf{W}}^{2}_{b}$)
$\left[\begin{array}[]{c}{\bf{X}}^{t+1}\\ {\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}{\bf{W}}_{a}&-{\bf{W}}_{a}\\ {\bf{I}}-{\bf{W}}_{a}&{\bf{W}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{X}}^{t}\\ {\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}{\bf{W}}_{a}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}{\bf{W}}_{b}^{2}{\bf{\Sigma}}_{1}^{t}+{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ {\bf{W}}_{b}{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\\ {\bf{I}}-\hat{{\bf{\Lambda}}}_{a}&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
@@
${\bf{H}}=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\\ {\bf{I}}-\hat{{\bf{\Lambda}}}_{a}&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})&-({\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}}))\\ \frac{1}{2\chi}({{\bf{I}}-\hat{{\bf{\Lambda}}}})&{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})\\ \end{array}\right],$
@@
$H_{i}=\left[\begin{array}[]{cc}1-\frac{1}{2\chi}(1-\lambda_{i})&-(1-\frac{1}{2\chi}(1-\lambda_{i}))\\ \frac{1}{2\chi}(1-\lambda_{i})&1-\frac{1}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
@@
Notice that $|\gamma_{(1,2),i}|<1$ when $-\nicefrac{{1}}{{3}}<\nu_{i}<1$, which holds under Assumption 1 since ${\bf{W}}_{a}\succ 0$, i.e., $0<\nu_{i}<1\ (i=2,\ldots,n)$. For $10<\nu_{i}<$ 1, the eigenvalues of $H_{i}$ are complex and distinct:
@@
Since the spectral radius of matrix is upper bounded by any of its norm and $10<\nu_{i}<$ 1, it holds that
@@
Using $\nu_{i}\geq 1-\frac{1}{2\chi}(1-\lambda_{n})$ and $1-\nu_{i}=\frac{1}{2\chi}(1-\lambda_{i})\geq\frac{1}{2\chi}(1-\lambda_{2})$, we have
$\|Q_{2,i}^{-1}\|^{2}\leq\frac{\chi}{(1-\frac{1}{2\chi}(1-\lambda_{n}))(1-\lambda_{2})}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}\ .$
@@
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
@@
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}.$
@@ V Proof of Lemma 3
$-\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big{\rangle}=-\frac{1}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}-\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$ $=\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\mathbb{F}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right].$
Since ${\bf{E}}^{t}=\frac{(\theta_{t}-1)}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\frac{1-p}{4\chi^{2}}\Big{(}\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\Big{)}\leq{\frac{2(1-p)}{\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+{\frac{2(1-p)}{\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{2(1-p)}{\chi^{2}}\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (50)
@@
$\mathbb{G}^{t}$ $={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]$
$={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\underbrace{\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]}_{{\bf{F}}^{t}}-\alpha\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}\\ \end{array}\right]}_{{\bf{C}}}{\bf{S}}^{t}$
@@
Note that ${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]$ and ${\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}$. It follows that

### v4

@@ VIII Proof Theorem 2
$0$ $=\nabla F({\bf{X}}^{\star})+{\bf{W}}_{b}{\bf{U}}^{\star},$ (45a)
$0$ $={\bf{W}}_{b}{\bf{Z}}^{\star},$ (45b)
@@
$\mathcal{E}_{\mathrm{s}}^{t}\triangleq\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]\text{ and }\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}<1.$
@@
Proof See Appendix XVII. $\Box$
With this error dynamics, similar as Lemma 3, we give the following descent inequalities.
@@
where $\vartheta_{\mathrm{s}}=\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}$, $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3p(1-p)(2+p^{2})}{\chi^{2}}$
@@
Proof See Appendix XVIII. $\Box$
@@
$\chi>\max\left\{1,\frac{36}{1-\lambda_{2}},\frac{72(1-p)}{1-\lambda_{2}}\right\},$ (49)
@@
Proof See Appendix XIX. $\Box$
IX Proof of Lemma 2
Proof Since $\widehat{{\bf{W}}}_{a}=\mathbf{I}-\frac{p}{2\chi}{\bf{W}}_{b}^{2}$, it follows from (2b), $\mathbf{I}-\mathbf{W}_{a}=\frac{1}{2\chi}\mathbf{W}_{b}^{2}$ and ${\bf{E}}^{t}=\frac{\theta_{t}-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$ that
@@ IX Proof of Lemma 2
$=(\mathbf{I}-\frac{1}{2\chi}{\bf{W}}_{b}^{2})\hat{\mathbf{Z}}^{t}+\frac{1-p}{2\chi}{\bf{W}}_{b}^{2}\hat{\mathbf{Z}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t}$
$=\widehat{{\bf{W}}}_{a}\hat{\mathbf{Z}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t}.$
Since $\beta=1$, it follows from (2b) that
${\bf{Y}}^{t+1}$ $=\mathbf{Y}^{t}+\hat{\mathbf{Z}}^{t}-\mathbf{X}^{t+1}$
$={\bf{Y}}^{t}+({\bf{I}}-\widehat{{\bf{W}}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}.$
By ${\bf{R}}^{t}={\bf{Y}}^{t}+\alpha\nabla F(\bar{{\bf{X}}}^{t})$, ${\bf{\Sigma}}_{2}^{t}=\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1})$, and ${\bf{E}}^{t}=\frac{\theta_{t}-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, we have
@@
$=({\bf{I}}-\widehat{{\bf{W}}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}.$
@@
${\bf{X}}^{t+1}$ $=\widehat{{\bf{W}}}_{a}\hat{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$
${\bf{R}}^{t+1}$ $={\bf{R}}^{t}+({\bf{I}}-\widehat{{\bf{W}}}_{a})\hat{{\bf{Z}}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}+{\bf{W}}_{b}{\bf{E}}^{t},$
which also can be rewritten as
$\left[\begin{array}[]{c}{\bf{X}}^{t+1}\\ {\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}\widehat{{\bf{W}}}_{a}&-\widehat{{\bf{W}}}_{a}\\ {\bf{I}}-\widehat{{\bf{W}}}_{a}&\widehat{{\bf{W}}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{X}}^{t}\\ {\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\widehat{{\bf{W}}}_{a}{\bf{\Sigma}}_{1}^{t}\\ \frac{p}{2\chi}{\bf{W}}_{b}^{2}{\bf{\Sigma}}_{1}^{t}+{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ {\bf{W}}_{b}{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\\ {\bf{I}}-\hat{{\bf{\Lambda}}}_{a}&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
@@
[…]
@@
$H_{i}=\left[\begin{array}[]{cc}1-\frac{p}{2\chi}(1-\lambda_{i})&-(1-\frac{p}{2\chi}(1-\lambda_{i}))\\ \frac{p}{2\chi}(1-\lambda_{i})&1-\frac{p}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{p}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
@@
Notice that $|\gamma_{(1,2),i}|<1$ when $\nicefrac{{1}}{{3}}<\nu_{i}<1$, which holds under Assumption 1 since $\widehat{{\bf{W}}}_{a}\succ 0$, i.e., $0<\nu_{i}<1\ (i=2,\ldots,n)$. For $0<\nu_{i}<1$, the eigenvalues of $H_{i}$ are complex and distinct:
@@
Since the spectral radius of matrix is upper bounded by any of its norm and $0<\nu_{i}<1$, it holds that
@@
Using $\nu_{i}\geq 1-\frac{p}{2\chi}(1-\lambda_{n})$ and $1-\nu_{i}=\frac{p}{2\chi}(1-\lambda_{i})\geq\frac{p}{2\chi}(1-\lambda_{2})$, we have
$\|Q_{2,i}^{-1}\|^{2}\leq\frac{\chi}{p(1-\frac{p}{2\chi}(1-\lambda_{n}))(1-\lambda_{2})}\leq\frac{2\chi}{p(1+\lambda_{n})(1-\lambda_{2})}\ .$
@@
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}<1.$
@@
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{p(1+\lambda_{n})(1-\lambda_{2})}.$
@@ X Proof of Lemma 3
$\big\langle\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big\rangle=-\frac{1}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}-\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]=$ $\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\mathbb{F}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[2\left\langle\mathbb{G}^{t},\mathbb{F}^{t}\right\rangle\;|\;\mathcal{F}^{t}\right].$
Since ${\bf{E}}^{t}=\frac{\theta_{t}-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
$\mathbb{E}\!\left[{\bf{E}}^{t}\;|\;\mathcal{F}^{t}\right]=p\ \frac{1-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}+(1-p)\ \frac{-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}=0.$
Noting that $\mathbb{F}^{t}=[-{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t};{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}]$, we have $\mathbb{E}\!\left[\left\langle\mathbb{G}^{t},\mathbb{F}^{t}\right\rangle\;|\;\mathcal{F}^{t}\right]=0$. On the other hand, we have
$\mathbb{E}\!\left[\|\mathbb{F}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$ $={\frac{p(1-p)^{2}}{2\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+{\frac{(1-p)p^{2}}{2\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}$
$\leq{{\frac{2p(1-p)}{\chi^{2}}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+{{\frac{2p(1-p)}{\chi^{2}}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+{\frac{2p(1-p)}{\chi^{2}}}\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (51)
@@
$\mathbb{G}^{t}$ $={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]$
$={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\underbrace{\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]}_{{\bf{F}}^{t}}-\alpha\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}\\ \end{array}\right]}_{{\bf{C}}}{\bf{S}}^{t}$
@@
Note that ${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]$ and ${\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}$. It follows that